OpenStudy (anonymous):

Solve by simplifying the problem. There are 18 students in Mr. Alvarez’s art class. Throughout the year, every student must pair up with every other student to complete a project. How many projects will be finished?

OpenStudy (anonymous):

This is a combination problem. n C 2 = [n (n-1)]\2 = (18 x 17) /2 = 153. you can count directly. Since the first student must pair up with the remaining 17, the second student must pair up with 16... or 17 + 16 + 15 + ... + 1 = (17x18)/2 = 153